OpenStudy (anonymous):

When you write the given number in scientific notation, what is the exponent in the power of 10? (Type only the value of the exponent.) 31,800,000=3.18 * 10^?

OpenStudy (anonymous):

just like : 100 = 1 * 10^2

OpenStudy (anonymous):

so its 2?

OpenStudy (anonymous):

31800000 = 318 * 100000 = 318 * 10^5 318/100 * 10^7 = 3.18 * 10^7

OpenStudy (anonymous):

ok so ^7

OpenStudy (anonymous):

No , it is 7 @rebecca1233 Check it again!

OpenStudy (anonymous):

Yeah it is 7

OpenStudy (anonymous):

please help some more

OpenStudy (anonymous):

7